You measure 41 turtles wroghrs and find they have a mean weight of 60 ounces. Assume the population standard deviation is 13.6 ounces. Based on this, construct a 95% confidence interval for the true population mean turtle weight.
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 60

Population standard deviation = \sigma = 13.6

Sample size = n = 41

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.96 * (13.6 / \sqrt41)

= 4.2

At 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

60 - 4.2 < \mu < 60 + 4.2

55.8 < \mu < 64.2

(55.8 , 64.2)